A pack of elephants are moving at 6 mph. One gets left behind. It chased at a rate of 10 mph and caught up in 5 minutes. Howlong in hours did it run to catch up? How far did the lone elephant run?
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)

A pack of elephants are moving at 6 mph. One gets left behind. It chased at a rate of 10 mph and caught up in 5 minutes. Howlong in hours did it run to catch up? How far did the lone elephant run?
Let the catchup distance be x miles
time to catchup 5 minutes= 5/60 hours = 1/12 hours
rate of lone elephant = 10mph
D= s*t
D= 1/12 * 10 =5/6 m

Answer by greenestamps(13200)



The solution from tutor @mananth assumes the pack of elephants stops to wait for the lone elephant to catch up, so that the lone elephant catches up to the pack at a rate of 10mph.

More likely, the pack kept moving at 6mph, so the rate at which the lone elephant catches up is 10-6 = 4mph.

ANSWERS:

The lone elephant ran for 5 minutes to catch up; 5 minutes is 1/12 of an hour.
The lone elephant ran for 1/12 of an hour at 10mph to catch up; the distance it ran to catch up was 10*(1/12) = 5/6 miles.